AIR-BROADENING AND SHIFT COEFFICIENTS AND LINE MIXING IN THE ν_{3} BAND OF ^{12}CH_{3}D

V. MALATHY DEVI, D. CHRIS BENNER, Department of Physics, The College of William and Mary, Box 8795, Williamsburg, VA 23187-8795; \underline{M. A. H. SMITH}, C. P. RINSLAND, Atmospheric Sciences, NASA Langley Research Center, Mail Stop 401A, Hampton, VA 23681-2199; and L. R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109.

A multispectrum nonlinear least squares fitting technique^{a} has been used to determine Lorentz air-broadening coefficients and air-induced shift coefficients for transitions in the ν_{3} fundamental band of ^{12}CH_{3}D in the spectral region between 1154 and 1430 cm^{−1}. Eleven high-resolution (0.005 cm^{−1}) room-temperature absorption spectra, recorded using the 1-m Fourier transform spectrometer (FTS) at the McMath-Pierce facility of the National Solar Observatory at Kitt Peak, were simultaneously analyzed. The data set included both low-pressure (1 to 3 Torr) spectra of 98% pure CH_{3}D and spectra of lean mixtures (≈ 1%) of CH_{3}D in dry air at total pressures from about 100 Torr to 400 Torr. Cell path lengths of 25 and 150 cm were used. Air-broadening coefficients were determined for for more than 360 ν_{3} transitions with rotational quantum numbers as high as J′′ = 17 and K′′ = 17. Air-induced shift coefficients were also determined for most of these transitions. The measured broadening coefficients range from 0.016 to 0.073 cm^{−1} atm^{−1} at 296K, and the shift coefficients range vary from about -0.0086 to +0.0058 cm^{−1} atm^{−1}. The majority of the shifts are negative, and the positive shifts often involve transitions with J′′ = K′′. The ^{Q}Q sub-band J′′ = K′′ transitions are also associated with the smallest broadening coefficients. Weak line mixing effects have been observed in a few high-J transitions with K′′ = 3, and we have determined offdiagonal relaxation matrix element coefficients for several A^{+}A^{−}(A1A2) split components. At low to medium values of J′′, the A^{+}A^{−} splittings are very small, and the two components are practically unresolved. Variations of the measured parameters with rotational quantum numbers and differences between the A and E symmetry species will be discussed. We will also compare our measurements with the values on the current HITRAN compilation^{b} and with other available measurements.

aD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith, and D. Atkins, JQSRT \underline{53}, 705-721 (1995).

bL. S. Rothman et al., JQSRT \underline{53}, 665-710 (1998).